2% of the trees in a plantation are known to have a certain disease.what is the probability that,in a sample of 250 trees.

(I) less than 1% are diseased


1
Expert's answer
2021-12-16T07:26:05-0500

"n=250 \\

p=0.02 \\

q = 1-p=1-0.02=0.98 \\

\hat{p} → N(\mu_{\hat{p}}, \sigma_{\hat{p}}) \\

\mu_{\hat{p}} = p = 0.02 \\

\sigma_{\hat{p}} = \sqrt{\frac{p \times q}{n}} \\

= \sigma{\frac{0.02 \times 0.98}{250}} = 0.0089"

P[less than 1% are diseased] = P(p<0.01)

"= P(Z< \frac{0.01-0.02}{0.0089}) \\

= P(Z< -1.12) \\

= 0.1314"

The required probability is 0.1314